Calculate percentage of purity of h2o2,0.2g of which has been reacted with 20 ml of decimolar kmno4 in acidic medium?

Grade:12th pass

1 Answers

Vikas TU
7 years ago
 2KMnO4 + H2O2 = 2KOH + MnO2 +O2 is the reacn.
Valence factor for KMnO4 => 7 – 4 = +3
Valence factor for H2O2 => +4
Hence moles for H2O2=>
20 * 0.1 * 3 = mmoles of H2O2 * 4
mmoles of H2O2 = 6/4 = 3/2 = 1.5
Weight of H2O2 = > 1.5*10^-3*34 = >  51*10^-3 gms.  => 0.051 gm.
percentage purity = > (0.2 – 0.051)*100/0.2 = > 74.5%